# Failure Modes of Maximum Entropy RLHF

### v1

@@ 4.1 Experimental Setup and Methodology
In our experiments, we train 1B and 2.8B parameter models from the Pythia suite (Biderman et al., 2023) using RLOO (Ahmadian et al., 2024) on the TL;DR dataset (Stiennon et al., 2022). For optimization, we follow the training recipe outlined in Huang et al. (2024), and implement our experiments using the TRL library (von Werra et al., 2020). We evaluate alignment to human preference using simulated win-rates with GPT-4o-mini (OpenAI et al., 2024) as the proxy evaluator, measured against reference summaries for TL;DR using greedy sampling unless stated otherwise.
Our model and dataset choices are guided by two main considerations. First, our aim is not to train state-of-the-art competitive models, but to study the methodological aspects of the approach. Second, computational constraints limit us from scaling to larger models. Nonetheless, this setup provides a well-suited testbed for exploring the questions we set out to investigate.
We adopt RLOO as a critic-free alternative to the standard RLHF pipeline, while still optimizing the same underlying objective. In the KL-constrained formulation, the reward is defined as
$r(x,y)=r_{\phi}(x,y)-\beta\Big(\log\pi_{\theta}(y|x)-\log\pi_{\text{ref}}(y|x)\Big),$ (17)
whereas in the length-normalized maximum-entropy variant, it takes the form
@@ 4.2 Results and Analysis
4.2.1 Online Maximum Entropy RLHF
Online Maximum Entropy RLHF with Pythia 1B.
To evaluate the effectiveness of Maximum Entropy RL, inspired by the success of SimPO, we trained the Pythia 1B model across a range of entropy coefficients and two learning rates: $1\times 10^{-6}$ and $1\times 10^{-7}$. The entropy coefficients were selected via a simple grid search, while the learning rates were motivated by prior findings, $1\times 10^{-6}$ being the setting where KL-constrained RLHF performs strongly, and $1\times 10^{-7}$ following the recommendation from SimPO. Our results, in Figure 1 , reveal that training with $1\times 10^{-6}$ consistently leads to overoptimization, regardless of the entropy coefficient. This suggests that entropy regularization alone is insufficient to constrain the model, which is ultimately unsurprising.
[…]
Online Maximum Entropy RLHF with with Pythia 2.8B.
To further validate our observations, we conducted experiments with Pythia 2.8B, a larger model from the same family. The only change in configuration was the learning rate, which we reduced to $1\times 10^{-7}$ and $5\times 10^{-8}$ to account for the increased model size; all other hyperparameters were kept as before , and results are given in Figure  2. Surprisingly, none of the trained variants were able to outperform the SFT baseline, as they all exhibited severe overoptimization. Notably, even the very small learning rate of $5\times 10^{-8}$ led to overoptimization, whereas the KL-constrained approach still achieved strong results with a higher learning rate of $1\times 10^{-6}$ under the same number of optimization steps. These findings highlight a clear failure case of Maximum Entropy RL as an online RLHF paradigm.
@@ 4.2.1 Online Maximum Entropy RLHF
Figure 3: KL divergence metrics and win rates for KL-Constrained and Maximum Entropy regularization methods across training steps. Top row shows KL divergence between current policy and reference policy ($KL(\pi_{t}||\pi_{\text{ref}})$) for 1B model (left) and 2.8B model (right). Bottom row shows KL divergence between consecutive policy iterations ($KL(\pi_{t}||\pi_{t-1})$)
KL Budget of the Optimization.
It is evident that Maximum Entropy RL is not sufficient to prevent overoptimization, and in fact, even RL without regularization can achieve comparable performance. Nevertheless, we observe overoptimization even at very small learning rates. To investigate this phenomenon, we track the KL divergence between the policy and the reference model during training in the Maximum Entropy setting. In addition, we consider two standard KL-constrained runs: one achieving a strong win rate and another that overoptimizes. The only difference between these two runs is the KL coefficient, which ultimately determines their outcomes. By comparing these cases, we aim to better understand KL behavior under the standard methodology and clarify what constitutes desirable optimization.
A key strength of a well-tuned KL regularizer is that, after some steady improvements, it ensures KL divergence grows only very slowly while the policy remains close to the reference model. This keeps optimization within safe regions. At the same time, KL regularization is highly sensitive, since even small changes in the KL coefficient can cause large shifts that ultimately lead to an overoptimized model. In addition, the appropriate KL weight is not universal and must be carefully tuned for each model, even when trained on the same dataset.
In our Maximum Entropy RL runs with the 1B model, in Figure  5, we find that the policy does not become overoptimized, but its KL grows in a nearly linear fashion, as expected. A similar pattern is observed with the 2.8B model; however, despite ending with almost identical KL values, the 2.8B model still collapses into an overoptimized state. This shows that the optimization budget in reference-free RL methods, whether Maximum Entropy RL or standard RL, is extremely fragile. Even with very small learning rates, models can still undergo significant KL updates that result in an overoptimized outcome.
The reason KL regularization is effective, despite only shaping the reward, is that it maintains a good KL divergence by penalizing out-of-distribution samples and dynamically dampening their effective reward. In contrast, Maximum Entropy RL cannot provide this safeguard because entropy correlates with overoptimization, which amplifies the issue rather than preventing it. Pure RL methods are even more vulnerable since they directly maximize the proxy reward, which inherently deteriorates once the policy drifts too far from the reference.
KL Update Magnitudes in Policy Optimization.
We observe that Pythia 2.8B exhibits high KL updates between consecutive policies even under very low learning rates. One could argue that in Maximum Entropy RL, such high KL updates arise not from the objective itself but from challenges in policy optimization. Since RLOO’s policy loss is implemented as PPO (with AC2 being a special case of PPO (Huang et al., 2022)), some may claim the issue is algorithmic, namely PPO’s difficulty in keeping ratios bounded (Wang et al., 2020), which ultimately destabilizes KL.
Our results in Figure  3, however, show that this is not purely algorithmic: in KL-constrained runs, PPO successfully maintains stable KL between successive policies, even in overoptimized settings where the KL constraint is relaxed. By contrast, in Maximum Entropy runs we consistently observe increasing KL drift, even in “good” runs, and this effect grows stronger during training despite using very low learning rates compared to KL-constrained runs. A plausible explanation is that unregularized reward optimization produces sharper gradients, which push the policy to change more aggressively. This is somewhat surprising in RLOO, since PPO’s average clipping ratio is quite low (unlike in standard RL), and the initial policy is already strong. Yet, in Maximum Entropy settings, we find a quadratic clipping behavior, suggesting that optimization drifts toward reward hacking regions that ignore regularization and focus solely on maximizing reward. To counteract this, we attempted to enforce stricter updates by reducing the PPO clipping parameter $\epsilon$ from the standard $0.2$ down to as small as $10^{-4}$ while using a learning rate of $5\times 10^{-8}$ in Pythia 2.8B experiments. However, this adjustment failed to induce greater pessimism: models still overoptimized, indicating that the problem is not resolved by clipping alone and may indeed be algorithmic.
Overall, our findings suggest that high KL is both an objective-driven and algorithmic phenomenon. KL-constrained runs remain stable (even when overoptimized), while Maximum Entropy runs show persistent KL escalation despite tighter clip ranges. This highlights that the optimization trajectory is strongly shaped by the choice of objective, even when using identical policy optimization techniques.
@@ 4.2.2 Minimum Entropy RL
Motivated by the link between maximum entropy and overoptimization and recent work showing entropy minimization can serve as an effective reward signal for LLM reasoning (Agarwal et al., 2025), we adopt an unconventional strategy: minimizing entropy to discourage excessively high-entropy which we expect to prevent overoptimization.
@@ 4.3 Offline Maximum Entropy RLHF (SimPO)
Even though Maximum Entropy fails to provide sufficient regularization to prevent overfitting, its closed-form solution, SimPO, proves to be both effective and performant. This effectiveness cannot be solely attributed to the use of a low learning rate, since one of the configurations for Llama 3 (Grattafiori et al., 2024) employs an even higher learning rate than DPO. Nevertheless, maintaining a low learning rate remains critical for controlling the KL, which is a crucial and universal requirement across all alignment algorithms (Gao et al., 2022; Rafailov et al., 2024a).
@@
One might argue that, since all samples are in-distribution, there is no need for an explicit OOD regularizer. However, as noted by Azar et al. (2023); Rafailov et al. (2024a), the reward model effectively drifts out-of-distribution during optimization, which leads to sub-epoch overoptimization. This highlights that the form of the reward model is critical, and that the mere presence of a reference model is insufficient. To mitigate this, Huang et al. (2025) propose replacing KL with $\chi^{2}$ regularization, thereby injecting pessimism directly into the reward model. They report that this approach maintains performance across multiple epochs, whereas DPO collapses after just one. However, we were unable to reproduce these results, leaving open the question of whether $\chi^{2}$ regularization truly implements pessimism. This makes the performance of SimPO particularly intriguing, not because it achieves pessimism, but because it demonstrates strong results even without relying on a reference model.
@@
In DPO, the pairwise reward can be written as $r(y_{w}|x)-r(y_{l}|x)=\beta\left(\log\frac{\pi(y_{w}|x)}{\pi(y_{l}|x)}-\log\frac{\pi_{\text{ref}}(y_{w}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$, where the second term reflects the contribution of the reference model. Because both $y_{w}$ and $y_{l}$ are sampled from the reference distribution, we expect this term to be negative but small, effectively acting as an adaptive regularizer. This parallels the role of a margin in SimPO, with the key distinction that SimPO uses a fixed margin rather than a reference-based one (Ahrabian et al., 2025).
@@
Figure 4: Batch average of $\log(\frac{\pi_{\text{ref}}(y_{w}|x)}{\pi_{\text{ref}}(y_{l}|x)})$ during DPO training.
This perspective suggests that offline methods might potentially reduce reliance on reference models by introducing target margins that could serve a similar function to reference contributions. To explore this possibility, we visualize the reference log probability margins $\log\left(\frac{\pi_{\text{ref}}(y_{w}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$ during DPO training with Pythia 1B, in Figure  4. Our observations suggest that these margins tend to fall within a relatively narrow range, which appears consistent with the fixed margins used in reference free methods like SimPO.
Some caveats are worth noting. High learning rates combined with large margins can drive aggressive optimization that maximizes separation, potentially leading to the reward overoptimization behaviors highlighted by Rafailov et al. (2024a). We observe extreme likelihood decreases, suggesting that the model places greater weight on out of distribution samples. Reference models may provide adaptive margins that guide optimization, with minimal contributions reducing overoptimization and larger margins focusing on harder examples. Yet, fixed margins risk forcing overoptimization under aggressive updates. These findings indicate that reference models are neither necessary nor sufficient to prevent overoptimization, as reward hacking in DAAs still emerges from overfitting to the reward objective, with cancellation effects limiting protection. We discuss these dynamics in more detail in the Appendix  B.
@@ 5 Conclusion
This work establishes a theoretical foundation for SimPO by connecting it to Maximum Entropy Reinforcement Learning with adaptive temperature scaling, while revealing a striking asymmetry between offline and online performance. Although SimPO excels in offline preference optimization, our empirical investigation shows that online Maximum Entropy RL suffers from instability and overoptimization, with entropy regularization paradoxically correlating with rather than preventing reward hacking. These findings highlight that reference-free approaches, while appealing for their simplicity, may face fundamental limitations in online training scenarios, and suggest that SimPO’s success stems from implicit stabilizing factors such as dataset constraints and target margins that approximate the regularization benefits of reference models.
Ethics Statement
This work focuses on the theoretical and empirical analysis of reinforcement learning objectives for aligning large language models. All experiments were conducted on publicly available preference datasets, and no personally identifiable or sensitive information was used. Our results are intended to improve the understanding of alignment methods and do not involve deployment of models in real-world settings. Nevertheless, as with all research on large language models, advances in alignment can have dual-use implications: while they may contribute to safer and more reliable AI systems, they could also lower barriers to developing more capable models that might be misused. We encourage responsible use and further investigation into the societal impacts of alignment research.
Reproducibiblity Statement
Our experiments are based on publicly available models (Pythia (Biderman et al., 2023)) and the TRL library (von Werra et al., 2020), with only minimal modifications. Because we rely primarily on standard, open-source components, our results are fully reproducible and can be replicated by other researchers.
The Use of Large Language Models
All text was initially drafted by the authors, after which Large Language Models were employed to refine phrasing and enhance clarity of expression.
@@ Appendix A Additional Related Work
Reward hacking (Skalse et al., 2025) is a long-standing problem in reinforcement learning (Sutton & Barto, 2018) where policies achieve high rewards but fail to meet the actual objective (Amodei et al., 2016; Hadfield-Menell et al., 2020; Pan et al., 2022). In language model alignment, this manifests as models learning to generate outputs that score highly on proxy metrics while being of poor actual quality. This overoptimization phenomenon was first systematically studied in traditional RLHF (Christiano et al., 2023; Stiennon et al., 2022; Gao et al., 2022; Ouyang et al., 2022), where optimizing imperfect proxy reward models leads to qualitatively worse outputs, including overly wordy responses and hallucinated information.

### v2

@@ 4.1 Experimental Setup and Methodology
We conduct experiments using 1B, 2.8B, and 6.9B parameter models from the Pythia suite (biderman2023pythiasuiteanalyzinglarge), trained with RLOO (ahmadian2024basicsrevisitingreinforcestyle) on the TL;DR summarization dataset (stiennon2022learningsummarizehumanfeedback). All experiments follow the RLHF training recipe described by huang2024nimplementationdetailsrlhf and are implemented using the TRL library (vonwerra2022trl). Model alignment is evaluated using simulated preference win rates computed with GPT-4o-mini (openai2024gpt4ocard) as a proxy evaluator, measured against reference summaries for TL;DR using greedy decoding unless otherwise stated. Our experimental protocol closely follows the setup of rafailov2024scalinglawsrewardmodel, enabling direct comparison with prior work.
Our aim is to study overoptimization phenomena in online RLHF, which are difficult to quantify reliably in more complex training setups involving larger models, diverse tasks, or heterogeneous reward signals. The TL;DR benchmark with Pythia models provides a controlled setting in which overoptimization can be measured consistently through preference win rates against fixed reference summaries, allowing systematic comparison across optimization objectives and model scales.
We adopt RLOO as a critic-free alternative to the standard RLHF pipeline while optimizing equivalent reward objectives. In the KL-constrained formulation, the reward is defined as
$r(x,y)=r_{\phi}(x,y)$
$-\beta\Big(\log\pi_{\theta}(y|x)-\log\pi_{\text{ref}}(y|x)\Big),$
whereas in the Maximum Entropy formulation, the reward takes the form
$r(x,y)=r_{\phi}(x,y)-\alpha\log\pi_{\theta}(y|x).$ (6)
When applying Maximum Entropy RL to sequence generation, the entropy term can be trivially increased by producing longer responses. Unlike KL-constrained objectives, which include an explicit reference policy that naturally counteracts such length effects, the reference-free formulation lacks an inherent mechanism to penalize verbosity. We therefore employ length normalization in the entropy term to ensure that the reward contribution is comparable across responses of different lengths and to prevent systematic length exploitation during optimization. This yields the reward
@@
We also explored extending our evaluation to additional preference datasets such as Anthropic-HH (bai2022traininghelpfulharmlessassistant). In this setting, we attempted to reproduce reported baselines using both the official DPO implementation and the TRL library. Across implementations, we consistently observed that DPO failed to outperform the SFT baseline, with performance often degrading during training. This behavior exhibited substantial variability across runs and implementations, making it difficult to establish stable and reproducible reference baselines for systematic analysis. We therefore focus our empirical study on the TL;DR benchmark, which provides more consistent training dynamics suitable for controlled investigation.
@@ 4.2 Results and Analysis
Online Maximum Entropy RLHF across model scales.
We evaluate Maximum Entropy RL in an online RLHF setting using Pythia models at three scales: 1B, 2.8B, and 6.9B parameters. Learning rates are chosen based on prior work, with $1\times 10^{-6}$ corresponding to a regime where KL-constrained RLHF performs well, and smaller learning rates following common practice in reference-free methods. Entropy coefficients are selected via grid search.
Across all model scales, Maximum Entropy RL is highly sensitive to the learning rate. At $1\times 10^{-6}$, all models rapidly enter overoptimized regimes regardless of the entropy coefficient. Lower learning rates improve stability, but do not fully eliminate this behavior.
For Pythia 1B, reducing the learning rate to $1\times 10^{-7}$ yields stable training and improved win rates relative to the SFT baseline (Figure 1). However, these gains are not driven by entropy regularization: similar performance is achieved even when the entropy coefficient is set to zero. Moreover, stable runs exhibit decaying entropy bonuses, while overoptimized runs show increasing entropy bonuses, suggesting that entropy correlates with reward hacking rather than preventing it.
Scaling behavior.
For Pythia 2.8B, we restrict experiments to learning rates of $1\times 10^{-7}$ and $5\times 10^{-8}$, as higher learning rates consistently caused rapid overoptimization. Despite these conservative settings, all Maximum Entropy variants overoptimize and fail to outperform the SFT baseline (Figure 2). In contrast, KL-constrained RLHF remains stable and effective at a higher learning rate of $1\times 10^{-6}$.
Interestingly, Pythia 6.9B does not exhibit behavior intermediate between smaller and larger models; instead, its training dynamics more closely resemble those of Pythia 1B than Pythia 2.8B. Training remains stable at a learning rate of $1\times 10^{-7}$ (Figure 11), but increasing the learning rate to $1\times 10^{-6}$ (Figure 10) leads to overoptimization, mirroring the instability observed in smaller models. However, KL-constrained RLHF remains stable for Pythia 6.9B while consuming only a small KL budget (Figure 12), indicating that effective KL budgets vary across model scales and are not reliably controlled by reference-free objectives.
KL budget and optimization dynamics.
To better understand these failures, we analyze KL divergence during training. Well-tuned KL-constrained RLHF exhibits an initial growth phase followed by slow KL increase, keeping the policy close to the reference model. This behavior is sensitive to the KL coefficient, but provides a reliable mechanism for controlling optimization.
In contrast, Maximum Entropy RL exhibits fragile KL behavior across model scales. For Pythia 1B and 6.9B, KL divergence grows approximately linearly without immediate collapse, whereas for Pythia 2.8B similar KL magnitudes correspond to severe overoptimization. This indicates that effective KL budgets are inherently model-dependent, and that reference-free objectives lack a mechanism to infer or enforce an appropriate optimization budget. As a result, their success is not predictable across model scales.
@@
Figure 3: KL divergence metrics for KL-constrained and Maximum Entropy RL across training. Top row shows KL divergence between the current policy and the reference policy ($KL(\pi_{t}\|\pi_{\text{ref}})$) for Pythia 1B (left) and 2.8B (right). Bottom row shows KL divergence between consecutive policy iterations ($KL(\pi_{t}\|\pi_{t-1})$).
KL update magnitudes and optimization stability.
To disentangle whether overoptimization arises from the objective or from policy optimization dynamics, we examine the KL divergence between consecutive policy updates. Since RLOO implements policy optimization using PPO, one might attribute instability in Maximum Entropy RL to PPO-specific issues, such as poor ratio control or clipping behavior.
However, Figures  3,  10,  11 show that this explanation is insufficient. In KL-constrained runs, PPO reliably maintains bounded KL updates between successive policies, even in regimes that eventually overoptimize when the KL coefficient is relaxed. In contrast, Maximum Entropy runs consistently exhibit increasing KL drift between updates, including in runs that appear well-behaved early in training. This effect intensifies over time despite using substantially lower learning rates than those required for stable KL-constrained optimization.
These observations suggest that the instability is not merely algorithmic, but objective-driven. Without an explicit reference policy anchoring the optimization, reward gradients become sharper and induce more aggressive parameter updates, leading to compounding KL drift. Tightening PPO clipping alone is insufficient to prevent this behavior: for Pythia 2.8B, we experimented with clipping ranges as small as $10^{-4}$, yet still observed increasing KL drift and eventual overoptimization. This indicates that reference-free objectives lack the structural safeguards needed to control update magnitudes in online RLHF.
@@ 4.2.1 Minimum Entropy RL
Motivated by the link between maximum entropy and overoptimization and recent work showing entropy minimization can serve as an effective reward signal for LLM reasoning (agarwal2025unreasonableeffectivenessentropyminimization), we adopt an unconventional strategy: minimizing entropy to discourage excessively high-entropy which we expect to prevent overoptimization.
@@ 4.3 Offline Maximum Entropy RLHF (SimPO)
Although Maximum Entropy RL fails to provide sufficient regularization in online settings, its closed-form offline solution, SimPO, is both effective and empirically competitive. This effectiveness cannot be attributed solely to conservative optimization: for example, one SimPO configuration for Llama 3 (grattafiori2024llama3herdmodels) uses a higher learning rate than DPO. Nevertheless, maintaining a sufficiently low learning rate remains critical for controlling the KL divergence, which appears to be a universal requirement across alignment methods (gao2022scalinglawsrewardmodel; rafailov2024scalinglawsrewardmodel).
@@
A common argument for offline stability is that all training samples remain in-distribution, removing the need for explicit out-of-distribution regularization. However, as noted by azar2023generaltheoreticalparadigmunderstand; rafailov2024scalinglawsrewardmodel, the reward model can effectively drift out of distribution during optimization, leading to sub-epoch overoptimization. This suggests that stability depends not only on data coverage, but also on the structure of the reward objective itself. To address this, huang2025correctingmythosklregularizationdirect propose replacing KL regularization with $\chi^{2}$ regularization to inject pessimism directly into the reward model, reporting improved stability across multiple epochs. We were unable to reproduce these results, leaving open the question of whether $\chi^{2}$ regularization reliably implements pessimism. In this context, SimPO’s performance is particularly intriguing, not because it enforces pessimism, but because it achieves strong results without relying on a reference model.
To better understand this behavior, consider the pairwise reward used in DPO:
$r(y_{w}|x)-r(y_{l}|x)$ $=\beta\Bigg(\log\frac{\pi(y_{w}|x)}{\pi(y_{l}|x)}$
$\quad-\log\frac{\pi_{\text{ref}}(y_{w}|x)}{\pi_{\text{ref}}(y_{l}|x)}\Bigg),$
@@
where the second term reflects the contribution of the reference model. Because both $y_{w}$ and $y_{l}$ are sampled from the reference distribution, this term is typically negative but small, acting as an adaptive regularizer. This behavior parallels the role of the target margin in SimPO, with the key distinction that SimPO employs a fixed margin rather than a reference-dependent one (ahrabian2025practicalanalysishumanalignment).
@@
Figure 4: Batch average of $\log\!\left(\frac{\pi_{\text{ref}}(y_{w}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$ during DPO training on Pythia 1B.
To examine this connection empirically, we visualize the reference log-probability margins
$\log\!\left(\frac{\pi_{\text{ref}}(y_{w}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$
during DPO training with Pythia 1B in Figure 4. We observe that these margins lie within a relatively narrow range, consistent with the fixed margins commonly used in reference-free methods such as SimPO. This suggests that, in offline settings, reference models may primarily serve to provide a bounded margin signal rather than strong distributional control.
Several caveats are worth noting. Large margins combined with high learning rates can drive aggressive separation, exacerbating reward overoptimization (rafailov2024scalinglawsrewardmodel). We observe extreme likelihood decreases indicative of such behavior, consistent with prior analyses of direct alignment algorithms. While reference models may implicitly adapt margins across samples—providing smaller corrections for easy examples and larger ones for harder cases—fixed margins lack this adaptivity and can amplify pathological updates under aggressive optimization. These findings suggest that reference models are neither necessary nor sufficient to prevent overoptimization in offline preference optimization. Instead, stability emerges from the interaction between learning rate, margin choice, and model capacity. We analyze these dynamics in more detail in Appendix B.
@@ 5 Conclusion
This work establishes a theoretical foundation for SimPO by connecting it to Maximum Entropy Reinforcement Learning, while revealing a striking asymmetry between offline and online performance. Although SimPO excels in offline preference optimization, our empirical investigation shows that online Maximum Entropy RL suffers from instability and overoptimization, with entropy regularization paradoxically correlating with rather than preventing reward hacking. These findings highlight that reference-free approaches, while appealing for their simplicity, may face fundamental limitations in online training scenarios, and suggest that SimPO’s success stems from implicit stabilizing factors such as dataset constraints and target margins that approximate the regularization benefits of reference models.
Acknowledgements
This work was funded by KUIS AI Center, Koç University.
@@ Appendix A Additional Related Work
Reward hacking (skalse2025definingcharacterizingrewardhacking) is a long-standing problem in reinforcement learning (Sutton1998) where policies achieve high rewards but fail to meet the actual objective (amodei2016concreteproblemsaisafety; hadfieldmenell2020inverserewarddesign; pan2022effectsrewardmisspecificationmapping). In language model alignment, this manifests as models learning to generate outputs that score highly on proxy metrics while being of poor actual quality. This overoptimization phenomenon was first systematically studied in traditional RLHF (christiano2023deepreinforcementlearninghuman; stiennon2022learningsummarizehumanfeedback; gao2022scalinglawsrewardmodel; ouyang2022traininglanguagemodelsfollow), where optimizing imperfect proxy reward models leads to qualitatively worse outputs, including overly wordy responses and hallucinated information.